OpenStudy (anonymous):

find the four arithmetic means between 12 and -13. i really need help

OpenStudy (anonymous):

I really don't know how to do this but can help you out :) You even have to solve

OpenStudy (anonymous):

First you have to subtract both the numbers 12-(-13)= ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What's the answer? :)

OpenStudy (anonymous):

i got 25

OpenStudy (anonymous):

Good correct :) If you want 4 arithmetic means you need to divide this by 5.

OpenStudy (anonymous):

okay i got 5:)

OpenStudy (anonymous):

In order to get 4 arithmetic means, just add 5, starting the smaller number -13

OpenStudy (anonymous):

-13+5=?

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

Yes now add 5 to -8

OpenStudy (anonymous):

-8+5=?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

Yes now add 5 to -3

OpenStudy (anonymous):

oh okay now i get it:) thank you:)

OpenStudy (anonymous):

What 4 numbers you are getting between 12 and -13 12, __ , __ , __ ,__ , -13

OpenStudy (anonymous):

7,2,-3,-8

OpenStudy (anonymous):

12,-8,-3,2,11,-13